OpenStudy (anonymous):

Find the possible values of x that satisfies the inequaty |x²-1|<=|x|

OpenStudy (anonymous):

answer is x<-17 -16-18>18+2x-18 -34>2x divide both sides by 2 and you get the answer

OpenStudy (anonymous):

Sorry?

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

Thats not correct.

OpenStudy (anonymous):

Do you know the absolute value laws?

OpenStudy (anonymous):

it is what is up there

OpenStudy (anonymous):

absolute value laws state that x = x if x>= 0 (If x is positive) and x = -x if x<0 (If x is negative)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but tell me how can i solve this problem? may I do first -x<= x²-1<=x

OpenStudy (anonymous):

than -x+1<=x²<=x+1 and now?

OpenStudy (anonymous):

no, I am going to write it out for you wait 2min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

constructing our laws: \[x^2-1 = x^2-1\] when \[x \ge 1\] \[x^2-1 = -(x^2-1)\] when \[x \lt 1\]

OpenStudy (anonymous):

\[x = x\] when \[x \ge 0\] and \[x = -x\] when \[x \lt 0\]

OpenStudy (anonymous):

do you understand so far? we are just setting up our laws.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

looking at our laws. We have 0 and 1. Can you see that? So we must now set up a NUMBER LINE with 0 and 1 so we can set up our inequalities. Number line: (Copy this and write it down) <------0-------1------>

OpenStudy (anonymous):

group up the inequalities in the laws for \[|x^2-1|\] and \[|x|\]

OpenStudy (anonymous):

oks and now?

OpenStudy (anonymous):



OpenStudy (anonymous):

the circle numbers are for the grouping, dont worry to much about that for now. Do you understand so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

i will try here. thanx

OpenStudy (anonymous):